OpenStudy (anonymous):

Use addition or substitution to find the value of y for this set of equations. 2x - 3y = -15 x = 4y

OpenStudy (anonymous):

substitute x=4y into the first equation. 2x - 3y = -15 2(4y) - 3y = -15

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Just substitute the value of x = 4y in place of x in first equation: 2(4y) - 3y = -15 8y - 3y = -15 Try to solve it and find y..